### v1

# Efficient classical algorithms for linear optical circuits

Abstract
We present efficient classical algorithms to approximate expectation values and probability amplitudes in linear optical circuits. Specifically, our classical algorithm efficiently approximates the expectation values of observables in linear optical circuits for arbitrary product input states within an additive error under a mild condition. This result suggests that certain applications of linear optical circuits relying on expectation value estimation, such as photonic variational algorithms, may face challenges in achieving quantum advantage. In addition, the (marginal) output probabilities of boson sampling with arbitrary product input states can be efficiently approximated using our algorithm, implying that boson sampling can be efficiently simulated if its output probability distribution is polynomially sparse. Moreover, our method generalizes Gurvits’s algorithm, originally designed to approximate the permanent, to also approximate the hafnian of complex symmetric matrices with an additive error. The algorithm also solves a molecular vibronic spectra problem for arbitrary product input states as precisely as boson samplers. Finally, our method extends to near-Clifford circuits, enabling the classical approximation of their expectation values of any observables and (marginal) output probabilities.
Introduction.— Quantum optical systems are promising platforms to realize quantum computational advantages. In particular, linear optical circuits are experimentally feasible and can be used to build universal quantum computers using the Knill-Laflamme-Milburn scheme [1]. Also, they are considered a promising platform to demonstrate the computational advantage in the near future since the seminal work on boson sampling [2]. The boson sampling proposal [2, 3, 4] motivated various experiments to demonstrate quantum advantages [5, 6, 7, 8, 9] while numerous classical algorithms have also been developed to challenge them, leading to significant advancements in both theoretical and experimental aspects of linear optics [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
With the rapid progress in boson sampling experiments, various proposals have emerged to leverage its advantage for solving practical problems [22, 23, 24, 25, 26, 27], though their quantum advantage remains unproven. A representative example is the molecular vibronic spectra problem [22], a chemistry problem solvable with Gaussian boson sampling [3, 4], which corresponds to computing grouped output probabilities of boson sampling. While many experiments have been conducted to demonstrate this proposal [28, 29, 30], Ref. [31] has recently introduced a classical algorithm that achieves comparable performance. Its key idea is that since Gaussian boson sampling computes quantities through sampling, the accuracy is limited by the finite sample size, which leaves room for a classical algorithm to achieve the same performance. This example illustrates that errors from finite sampling must be properly accounted for in order to rigorously evaluate any potential quantum advantage in computing physical quantities via sampling.
Another notable proposal for applying boson sampling to practical problems is variational quantum computing [32, 33, 34, 35, 36]. Similarly, boson samplers play a key role in these problems by computing physical quantities through sampling, meaning that the results are again inherently approximations with additive errors due to finite sampling. This potentially allows classical algorithms to match their performance. Therefore, understanding the complexity of approximating expectation values of physical quantities in linear optical circuits is crucial for assessing the potential quantum advantage of boson sampling in practical applications.
Figure 1: Summary of various classical algorithms introduced in the main text. (a) A schematic of the general expectation value of an $M$-mode linear optical circuit (Theorem 1) or an $M$-qubit near-Clifford circuit (Theorem 3). If $\hat{O}_{i}$ are photon-number projectors $|m_{i}\rangle\langle m_{i}|_{i}$, the expectation values represent marginal output probabilities of a boson sampling circuit, resulting in an efficient simulation of a sparse boson sampling (Corollary 1). (b) A probability amplitude of a linear optical circuit stands for generalized Gurvits’s algorithm (Theorem 2). When $|\phi_{i}\rangle$’s are single-photon Fock states and $|\psi_{i}\rangle$’s are squeezed vacuum states, the amplitude is proportional to the hafnian of a general complex symmetric matrix (Corollary 2). A probability amplitude is reduced to a phase shifter expectation value, Eq. (8), with $|\phi_{i}\rangle=|\psi_{i}\rangle$ and $\hat{U}\rightarrow\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ (Corollary 3).
In this work, we present a classical algorithm that efficiently approximates the expectation value of operators with respect to the output states of a linear optical circuit applied to an arbitrary product input state, provided that the norm of the operator of interest is well-bounded. Our algorithm suggests that certain applications proposed to leverage boson sampling’s quantum advantage can be dequantized if the purpose of using boson sampling is to approximate physical quantities, such as cost functions in variational quantum algorithms [34, 35]. In addition, our classical algorithm enables an efficient approximation of the (marginal) output probabilities of boson sampling circuits for arbitrary input states. Thus, it proves that any boson sampling can be classically simulated if the output probability distribution is polynomially sparse [37], extending previous results that applied only to restricted cases [38, 39, 40].
Furthermore, we present another classical algorithm that fully generalizes Gurvits’s algorithm [41, 2], which efficiently approximates probability amplitudes of linear optical circuits. It provides a quantum-inspired classical algorithm to approximate matrix functions, such as hafnian [42], by mapping the matrix into a Gaussian boson sampling’s output amplitude [3, 43]. Our algorithm can also efficiently approximate the expectation value of a phase shifter. The latter enables us to solve a generalized molecular vibronic spectra problem with arbitrary input states, a candidate for achieving a practical quantum advantage of boson sampling [31]. Additionally, it can be used for approximating binned probabilities of linear optical circuits with arbitrary input states, which have been proposed for verifying boson sampling experiments [44, 45, 46, 47].
Finally, using a similar procedure, we extend our algorithm to near-Clifford circuits [48, 49] and show that we can efficiently estimate the expectation values of observables and any (marginal) output probabilities for these circuits within an additive error.
Efficient approximation of expectation values.— Let $|\psi\rangle=\otimes_{i=1}^{M}|\psi_{i}\rangle$ be a product input state in an $M$-mode(-qubit) system, and $\hat{U}$ be a quantum circuit that is applied to the input state. For the output state of the circuit, $\hat{U}|\psi\rangle$, we consider the expectation value of a product operator $\hat{O}=\otimes_{i=1}^{M}\hat{O}_{i}$ (see Fig. 1(a)):
$\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle.$ (1)
Throughout this work, input states $|\psi\rangle$ and operators $\hat{O}$ are assumed to be in a product form unless stated otherwise, which is practically relevant in various applications. When an operator acts nontrivially only on a subsystem $A$ and trivially on the rest of the system $B$, we write the operator as $\hat{O}=\hat{O}_{A}\otimes\hat{\mathbb{1}}_{B}=\bigotimes_{i\in A}\hat{O}_{i}\otimes\bigotimes_{i\in B}\hat{\mathbb{1}}_{i}$ and define the reduced density matrix $\hat{\rho}_{A}\equiv\mbox{Tr}_{B}[\hat{U}|\psi\rangle\langle\psi|\hat{U}^{\dagger}]$.
We first focus on a linear optical circuit $\hat{U}$. Our first main result is a classical algorithm that efficiently approximates the expectation values of operators in a linear optical circuit:
Theorem 1 (Expectation value approximation in linear optical circuit).
Consider an $M$-mode linear optical circuit $\hat{U}$ and $K$ operators ${\{\hat{O}^{(k)}\}_{k=1}^{K}}$. The expectation values $\langle\psi|\hat{U}^{\dagger}\hat{O}^{(k)}\hat{U}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$.
Proof sketch.
(See SM Sec. S2 [50] for the full proof.) Assume that an operator $\hat{O}$ acts nontrivially on an $l$-mode subsystem $A$. Using the displacement operator expansion of the operator, $\hat{O}_{A}=\pi^{-l}\int d^{2l}\alpha\chi_{\hat{O}_{A}}(\bm{\alpha})\hat{D}^{\dagger}(\bm{\alpha})$ [51, 52, 53, 50], where $\hat{D}(\bm{\alpha})$ is a displacement operator and $\chi_{\hat{Q}}(\bm{\alpha})\equiv\mbox{Tr}[\hat{Q}\hat{D}(\bm{\alpha})]$ is the characteristic function of the operator $\hat{Q}$, we rewrite the expectation value as
$\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ $=\frac{1}{\pi^{l}}\int d^{2l}\alpha\chi_{\hat{O}_{A}}(\bm{\alpha})\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})$ (2)
$=\int d^{2l}\alpha\frac{|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}}{\pi^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})}{\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})},$ (3)
where $q(\bm{\alpha})\equiv|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}/(\pi^{l}\|\hat{O}_{A}\|_{2}^{2})$ is normalized to 1 and thus a proper probability distribution of $\bm{\alpha}$. Thus, the expression implies that by sampling $\bm{\alpha}$ from $q(\bm{\alpha})$ and averaging the random variable $X(\bm{\alpha})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})/\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})$ over the samples, we obtain an estimate of the expectation value.
We then show that the procedure can be efficiently performed and that the median-of-means estimator [54] yields an estimate with additive error $\epsilon$ with probability $1-\delta$ when $O(\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}\log\delta^{-1}/\epsilon^{2})$ number of samples is given, by showing that the variance of the random variable $X(\bm{\alpha})$ is upper-bounded by $\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}$. Finally, due to the union bound, we can approximate $K$ observables using $O(\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$ samples. Since it takes $O(M^{2})$ time for computing random variable $X(\bm{\alpha})$, the total running time is $O(M^{2}\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$.
Therefore, we achieve an efficient additive estimation of the expectation value in polynomial time when $\lVert\hat{O}\rVert^{2}_{2}=O(\text{poly}(M))$, a condition that holds in many interesting cases, as elaborated below.
Theorem 1 has various intriguing implications. First, any applications of linear optical circuits, such as boson sampling, that rely on estimating a physical observable $\hat{O}$ can be dequantized if the target observable’s norm is properly bounded; thus, this establishes a boundary for a classically simulable regime when it comes to approximating expectation values. Furthermore, when $\hat{O}$ is an orthogonal projector, its expectation value directly corresponds to the associated output probability. Therefore, if the rank of the orthogonal projector is well-bounded, the output probability of a linear optical circuit can be efficiently approximated. A notable example is a photon-number projector in boson sampling [2]. In this case, Theorem 1 implies that the (marginal) output probabilities of boson sampling with arbitrary input states can be efficiently approximated. To the best of our knowledge, such a result has only been known for the Fock state input [2, 45, 31]. Remarkably, this finding also enables an efficient simulation of boson sampling with arbitrary input states if polynomially many outcomes dominate the output probabilities [37, 55].
Corollary 1 (Sparse boson sampling).
Boson sampling with an arbitrary input state can be efficiently simulated if its output probability distribution is polynomially sparse.
We emphasize that Corollary 1 is valid for a general case regardless of phase space negativity [56, 40] or stellar rank [57] of the input state, resources for quantum advantages.
Another interesting example is recently proposed photonic variational quantum algorithms [34, 35, 36], whose cost function is written in terms of the expectation value as
$f(\theta)=\langle\psi|\hat{V}_{2}^{\dagger}e^{-i\theta\hat{n}_{k}}\hat{V}_{1}^{\dagger}\hat{O}\hat{V}_{1}e^{i\theta\hat{n}_{k}}\hat{V}_{2}|\psi\rangle,$ (4)
where $\hat{V}_{1}$ and $\hat{V}_{2}$ are linear optical circuits and $\hat{n}_{k}$ is the number operator on the $k$th mode. The main subroutine for obtaining the gradient is the estimation of $f(\theta)$’s for several values of $\theta$ by running the corresponding linear optical circuits. Since $\hat{V}_{1}e^{i\theta\hat{n}_{k}}\hat{V}_{2}$ is also a linear optical circuit, the estimation of the cost function can be replaced by the classical algorithm of Theorem 1 with $\hat{U}=\hat{V}_{1}e^{i\theta\hat{n}_{k}}\hat{V}_{2}$. Thus, our classical algorithm can replace the quantum algorithms when the norm of the operator $\hat{O}$ is in a product form and bounded.
Generalized Gurvits’s algorithm.— We now present another classical algorithm approximating a probability amplitude $\langle\phi|\hat{U}|\psi\rangle$ (See Fig. 1 (b)). Note that when $|\phi\rangle$ and $|\psi\rangle$ are single-photon Fock states, the amplitude becomes a permanent [2], which can be efficiently approximated by Gurvits’s algorithm [41]. Hence, we call our classical algorithm a generalized Gurvits’s algorithm, which fully extends the original Gurvits’s algorithm in the context of quantum optics:
Theorem 2 (Generalized Gurvits’s algorithm).
The probability amplitude $\langle\phi|\hat{U}|\psi\rangle$ of a linear optical circuit $\hat{U}$ with arbitrary states $|\psi\rangle$ and $|\phi\rangle$ can be efficiently approximated.
Proof Sketch.
We rewrite the amplitude by exploiting the completeness relation $\hat{\mathbb{1}}_{1:M}=\int d^{2M}\alpha|\bm{\alpha}\rangle\langle\bm{\alpha}|/\pi^{M}$:
$\langle\phi|\hat{U}|\psi\rangle$ $=\frac{1}{\pi^{M}}\int d^{2M}\alpha\langle\phi|\hat{U}|\bm{\alpha}\rangle\langle\bm{\alpha}|\psi\rangle$ (5)
$=\int d^{2M}\alpha\frac{\langle\phi|\hat{U}|\bm{\alpha}\rangle}{\langle\bm{\alpha}|\psi\rangle^{*}}\frac{|\langle\bm{\alpha}|\psi\rangle|^{2}}{\pi^{M}},$ (6)
where $|\bm{\alpha}\rangle$ represents a coherent state. Thus, if we sample $\bm{\alpha}$ from $|\langle\bm{\alpha}|\psi\rangle|^{2}/\pi^{M}$, which is a proper probability distribution (Husimi Q-distribution of the state $|\psi\rangle$ [58]) and easy to sample from, and average $\langle\phi|\hat{U}|\bm{\alpha}\rangle/\langle\bm{\alpha}|\psi\rangle^{*}$ over the samples, we obtain an estimate of the amplitude. We prove that this procedure is efficient and gives an estimate with $\epsilon$ additive error with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$ in SM Sec. S3 [50]. ∎
Note that the original Gurvits’s algorithm applies only when $|\psi\rangle=|\bm{m}\rangle$, $|\phi\rangle=|\bm{n}\rangle$ with $\bm{n},\bm{m}\in\{0,1\}^{M}$, and its variants [59, 31] apply when $\bm{m}\in\{0,1\}^{M}$ or $\bm{n}=\bm{m}$, corresponding to approximating the permanent of matrices with repeated rows and columns in a certain pattern. Not only does our algorithm extend to computing the permanent of matrices with repeated rows and columns in a general pattern, but it also introduces a new classical approach to approximating the permanent using a quantum-optics-inspired method.
Our algorithm can also approximate the hafnian [42], a generalization of the permanent, by using the fact that the hafnian of any $M\times M$ complex symmetric matrix $R$ ($M$ is even) with the largest singular value $\lambda_{\text{max}}<1$ can be expressed as a probability amplitude of an $M$-mode Gaussian boson sampling circuit [3, 43]:
$\text{haf}(R)=\frac{1}{Z^{1/2}}\langle\bm{1}|\hat{U}|\bm{r}\rangle,$ (7)
where $|\bm{r}\rangle=\otimes_{i=1}^{M}|r_{i}\rangle$ denotes a product of squeezed vacuum states $|r_{i}\rangle$ of squeezing parameter $r_{i}$, $|\bm{1}\rangle=\otimes_{i=1}^{M}|1\rangle$ is $M$ single-photon states, and $Z\equiv\prod_{i=1}^{M}\cosh{r_{i}}=\prod_{i=1}^{M}(1-\lambda_{i}^{2})^{-\frac{1}{2}}$ with $R$’s singular values $\lambda_{i}$. Until now, efficient additive approximation algorithms have been known only for restricted cases [40, 31]. Notably, our generalized Gurvits’s algorithm leads to an additive approximation algorithm for the hafnian without any constraints, which can be seen by substituting $|\psi\rangle=|\bm{r}\rangle$, $|\phi\rangle=|\bm{1}\rangle$ in Theorem 2:
Corollary 2 (Hafnian approximation).
The hafnian of an $M\times M$ complex symmetric matrix $R$ can be approximated within additive error $\epsilon\norm{R}^{M/2}$ with probability $1-\delta$ in $O(M^{3}\log\delta^{-1}/\epsilon^{2})$ time.
A detailed proof is provided in SM Sec. S4 [50]. By a similar method, the loop hafnian of an arbitrary complex symmetric matrix [43] also can be efficiently approximated.
It is worth emphasizing that Theorem 2 can also be used to approximate the overlap of two quantum states generated by linear optical circuits. Such a quantity has attracted attention for machine learning applications of boson sampling, such as kernel method [60, 32, 61]. More specifically, boson samplers are used to estimate the overlap of two quantum states corresponding to two different data vectors $\bm{x},\bm{y}$ as $|\langle\psi|\hat{U}^{\dagger}(\bm{x})\hat{U}(\bm{y})|\psi\rangle|^{2}$, where $\hat{U}(\bm{x})$ corresponds to the linear optical circuits whose configuration depend on the data vector $\bm{x}$. Thus, our classical method can replace boson samplers for this purpose.
Phase shifter.— We now consider a problem of approximating the expectation value of a phase shifter $e^{i\hat{\bm{n}}\cdot\bm{\phi}}$ of a linear optical circuit [62]
$\langle\psi|\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}|\psi\rangle,$ (8)
where $\hat{\bm{n}}\equiv(\hat{n}_{1},\dots,\hat{n}_{M})$ and $\bm{\phi}\in\mathbb{R}^{M}$ represents the phase vector. The expectation value of a phase shifter has various applications such as the molecular vibronic spectra problem [22, 31] and approximating binned probability, which has been proposed for verifying boson sampling experiments [44, 45, 46, 47].
In particular, the molecular vibronic spectra problem [22] has been considered as a potential application of boson sampling, which is essentially equivalent to computing grouped output probabilities $p(\bm{m})$ of boson sampling (See SM Sec. S4 [50] for more details):
$G(\Omega)\equiv\sum_{\bm{m}=\bm{0}}^{\infty}p(\bm{m})\delta(\Omega-\bm{\omega}\cdot\bm{m}),$ (9)
where $\bm{\omega}\in\mathbb{Z}^{M}_{\geq 0}$ and $\Omega=\{0,\dots,\Omega_{\text{max}}\}$. While a boson sampler can efficiently approximate Eq. (9) by sampling and counting the frequency of each group indexed by $\Omega$, classical approaches seem unlikely to be promising not just because the output probability $p(\bm{m})$ is #P-hard to compute but also because the number of outcomes $\bm{m}$ corresponding to each group $\Omega$ is generally exponentially large. In Ref. [31], however, the authors observe that the Fourier components of Eq. (9) are expressed as the expectation value of a phase shifter as Eq. (8). Therefore, if Eq. (8) is additively approximated, the grouped probabilities Eq. (9) are reproduced within an additive error through the inverse Fourier transform [50].
While Ref. [31] proves that this is the case when $|\psi\rangle$ is a product Gaussian state or a Fock state, leading to dequantization, for more general input states, like a product squeezed Fock state, a quantum advantage may still exist since no classical algorithm is known yet. Notably, by setting $|\phi\rangle=|\psi\rangle$ and $\hat{U}\to\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ in Theorem 2, our classical algorithm can approximate Eq. (8) for any product input state:
Corollary 3 (Phase shifter expectation value).
Phase shifter expectation value of a linear optical circuit can be efficiently approximated for an arbitrary product input state.
Therefore, our algorithm enables efficient approximation of the grouped probabilities in Eq. (9) for arbitrary input states and thus solves a generalized molecular vibronic spectra problem for arbitrary input states, an open problem posed in Ref. [31].
Extending to qubit circuits.— Finally, we show that our classical algorithm in Theorem 1 can be extended to near-Clifford circuits $\hat{U}_{\text{NC}}$ [48, 49], i.e., Clifford circuit $\hat{U}_{\text{C}}$+logarithmic number of $T$ gates, with arbitrary input states. Crucial properties used in Theorem 1 are that a displacement operator is transformed into another displacement operator through linear optical circuits and that any operator can be expanded in terms of displacement operators. Similar properties hold for Clifford circuits where any operators can be decomposed using Pauli operators and a Pauli operator transforms another Pauli operator under Clifford circuits. Thus, we obtain a similar result:
Theorem 3 (Expectation value approximation in near-Clifford circuit).
Consider an $M$-qubit near-Clifford circuit $\hat{U}_{\text{NC}}$ with depth $O(\text{poly}(M))$ and $K$ operators ${\{\hat{O}^{(k)}\}_{k=1}^{K}}$. The expectation values $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}^{(k)}\hat{U}_{\text{NC}}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(\text{poly}(M)\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$.
Proof Sketch.
The proof is similar to that of Theorem 1 and provided in SM Sec. S6 [50]. Let us first focus on Clifford circuits. Using the Pauli operator expansion, $\hat{O}_{A}=2^{-l}\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\chi_{\hat{O}_{A}}(\bm{a})\hat{P}_{\bm{a}}$, where $\chi_{\hat{Q}}(\bm{a})\equiv\mbox{Tr}[\hat{Q}\hat{P}_{\bm{a}}]$ and $\hat{P}_{\bm{a}}\in\{I,X,Y,Z\}^{l}$ is a Pauli operator, the expectation value of $\hat{O}=\hat{O}_{A}\otimes\hat{\mathbb{1}}_{B}$ can be written as
$\langle\psi|\hat{U}^{\dagger}_{\text{C}}\hat{O}\hat{U}_{\text{C}}|\psi\rangle=\frac{1}{2^{l}}\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\chi_{\hat{O}_{A}}(\bm{a})\chi_{\hat{\rho}_{A}}(\bm{a})$ (10)
$=\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\frac{|\chi_{\hat{O}_{A}}(\bm{a})|^{2}}{2^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})}{\chi^{*}_{\hat{O}_{A}}(\bm{a})}.$ (11)
Thus, by sampling $\bm{a}$ from $q(\bm{a})\equiv|\chi_{\hat{O}_{A}}(\bm{a})|^{2}/(2^{l}\|\hat{O}_{A}\|_{2}^{2})$ and averaging the random variable $X(\bm{a})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})/\chi^{*}_{\hat{O}_{A}}(\bm{a})$ over the samples, we obtain an estimate of the expectation value. By showing that the variance of the random variable $X(\bm{a})$ is bounded by $\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}$, the median-of-means estimator together with the union bound gives the desired error presented in the theorem.
Furthermore, our method is still valid for the additional logarithmic number of $T$-gates because a single Pauli operator is transformed into the sum of at most two Pauli operators after passing through each $T$-gate, that is, $\hat{T}^{\dagger}\hat{P}_{0}\hat{T}=\hat{P}_{0},~{}\hat{T}^{\dagger}\hat{P}_{1}\hat{T}=(\hat{P}_{1}+\hat{P}_{2})/\sqrt{2},~{}\hat{T}^{\dagger}\hat{P}_{2}\hat{T}=(\hat{P}_{2}-\hat{P}_{1})/\sqrt{2},~{}\hat{T}^{\dagger}\hat{P}_{3}\hat{T}=\hat{P}_{3}$. If there are $t$ of $T$-gates in a near-Clifford circuit $\hat{U}_{\text{NC}}$, we have at most $M2^{t}$ Pauli operators to track the transformations. Therefore, if $t=O(\log M)$, we can handle all the relevant Pauli operators in $O(\text{poly}(M))$ time. A detailed proof is given in SM Sec. S6 [50]. ∎
Theorem 3 implies that the expectation value of observables $\hat{O}$ with $\lVert\hat{O}\rVert_{2}^{2}=O(\text{poly}(M))$ and marginal probabilities of a near-Clifford circuit with an arbitrary input state can be efficiently approximated. Thus, for approximating the expectation values of near-Clifford circuits, the magic of input states does not increase the complexity [63]. Hence, our result is a generalization of previous results: the expectation values of a near-Clifford circuit for $|\psi\rangle=|0^{M}\rangle$ [48, 49] and (marginal) output probabilities of a Clifford circuit [64, 55].
Discussion.— We presented an efficient classical algorithm for estimating expectation values of observables in linear optical circuits. Using this, we showed that one can efficiently approximate (marginal) output probabilities of boson sampling with arbitrary input states, which implies that the simulation of sparse boson sampling with arbitrary input states is easy. We also generalized Gurvits’s algorithm to more general amplitudes that appear in linear optical circuits, which allows us to efficiently approximate a hafnian. Furthermore, using this classical algorithm, we solved an open problem in Ref. [31], namely, a generalized molecular vibronic spectra problem for arbitrary input states.
A natural open question is to extend our method to more complicated optical circuits with nonlinear effects such as nonlinear gates or postselection. Since such cases ultimately enable universal quantum computation [65, 1], we do not expect our method to work for the most general cases. Nonetheless, it suggests a phase transition from easiness to hardness when introducing nonlinear effects; thus, an interesting open question is to determine the exact transition point.
Also, we have focused on computing grouped probabilities of boson sampling with a certain pattern in the molecular vibronic spectra problem, which is defined by the linear relation $\Omega-\bm{\omega}\cdot\bm{m}$ in Eq. (9). However, our method does not straightforwardly generalize to a more complicated pattern of groups, which still opens a possibility of quantum advantage. Thus, the complexity of approximating general grouped boson sampling output probabilities is another open question.
Acknowledgements.
Acknowledgments.— This work was supported by Quantum Technology R&D Leading Program (Quantum Computing) (RS-2024-00431768) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT (MSIT)).