### v3

# Classical algorithms for estimating expectation values in linear-optical circuits

Abstract
We present a classical algorithm for approximating the expectation values of observables in linear-optical circuits with arbitrary product input states, achieving additive-error accuracy. This result indicates that current applications of photonic systems aimed at demonstrating practical quantum supremacy through expectation value estimation, such as photonic variational algorithms, may face challenges in attaining the computational advantage. It also implies the output probabilities of boson sampling with arbitrary product input states can be efficiently approximated by our method, resulting that boson sampling becomes efficiently simulable when its output probability distribution is polynomially sparse. We also develop an efficient classical algorithm for estimating transition amplitudes of arbitrary product states in linear-optical circuits. This provides additive-error approximation algorithms for matrix functions associated with linear-optical circuits, such as the (loop-)hafnian, which are of independent interest. As an application, it solves the generalized molecular vibronic spectra problem (Oh et al., 2024), previously suggested as a candidate for practical quantum advantage. Finally, we extend our framework to near-Clifford circuits, enabling classical approximation of their expectation values.
Introduction.— Linear-optical system has been considered as a promising platform for universal quantum computing [1] and also for demonstrating the quantum computational advantage in the near future since the seminal work on boson sampling [2]. The boson sampling proposal [2, 3, 4] motivated various experiments to demonstrate the quantum advantage [5, 6, 7, 8, 9] while numerous classical algorithms have also been developed to challenge them, leading to significant advancements in understanding the computational complexity of linear-optical circuits [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
Recently, the focus has shifted from demonstrating hardness through sampling to identifying practically motivated tasks that reflect a practical form of quantum advantage. A common viewpoint is that near-term quantum devices aim to estimate expectation values of observables with controlled additive accuracy, a task believed to be classically hard [24, 25, 26, 27, 28, 29, 30]. Within linear-optics, several proposals adopt this estimation-based framework [31, 32, 33, 34, 35], yet no provable advantage has been established. A notable case is molecular vibronic spectra (MVS): while Gaussian boson sampling enables an efficient photonic procedure [31, 3, 4, 36, 37, 38], recent work shows that a classical algorithm achieves comparable performance [39], because finite sampling enforces an additive-error tolerance, ruling out an advantage for MVS in this setting.
The essential issue, therefore, is precision: quantum devices estimate expectation values from finite samples, making additive-error approximations intrinsic to measurement (see SM Sec. S2 [40]). This immediately opens room for classical counterparts operating under the same additive tolerance. Accordingly, a rigorous assessment of practical quantum advantage requires identifying the classical limit of efficient estimability for the same observables within an additive-error tolerance. Although additive-error estimators for expectation values have been analyzed in other architectures [24, 41, 26, 42, 43, 44], to the best of our knowledge no such provable algorithm existed for linear-optical circuits.
Figure 1: Summary of classical algorithms introduced in the main text. (a) A schematic of the general expectation value of an $M$-mode linear-optical circuit (Theorem 1). If $\hat{O}_{i}$ are photon-number projectors $|m_{i}\rangle\langle m_{i}|_{i}$, the expectation values represent marginal output probabilities of a boson sampling circuit, resulting in an efficient simulation of a sparse boson sampling (Corollary 1). (b) A transition amplitude of a linear-optical circuit (Theorem 2). When $|\phi_{i}\rangle$’s are single-photon Fock states and $|\psi_{i}\rangle$’s are squeezed vacuum states, the amplitude is proportional to the hafnian of a general complex symmetric matrix (Corollary 2). A transition amplitude is reduced to a phase shifter expectation value, Eq. (6), with $|\phi_{i}\rangle=|\psi_{i}\rangle$ and $\hat{U}\rightarrow\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ (Corollary 3).
In this work, we present a classical algorithm for approximating expectation values in linear-optical circuits within additive-error. When the observable $\hat{O}=\hat{O}_{A}\otimes\mathbb{1}_{B}$} has polynomially bounded Hilbert–Schmidt norm, $\|\hat{O}_{A}\|_{2}^{2}=O(\mathrm{poly}(M))$, the algorithm is efficient for arbitrary product input states. This dequantizes a broad class of estimation tasks relevant to NISQ photonic computing that align with the framework of practical quantum advantage, including photonic variational quantum algorithms (VQAs) where each term of the target Hamiltonian is a local observable [45, 46, 47, 48, 49, 50]. The same framework efficiently approximates (marginal) output probabilities of boson-sampling circuits with arbitrary product inputs. Hence, any boson sampling is classically simulable if its output distribution is polynomially sparse [51], generalizing earlier restricted results [52, 53, 54]. Consequently, in the polynomially sparse regime, peaks of inverse-polynomial probability can be classically identified, tempering peak-based proposals of verifiable quantum advantage based on linear optics [55, 56, 57, 51].
We further develop an efficient quantum-optics-inspired algorithm for transition amplitudes in linear-optical circuits with arbitrary product inputs. While the single-photon Fock case involves permanents, classically estimable via Gurvits’ algorithm [58, 2], our analysis follows a distinct route with comparable guarantees and extends to quantum-inspired estimators for other matrix functions, including the (loop-)hafnian [59, 3, 60]. It also provides a complete classical solution to the generalized MVS problem [39] and to efficient estimation of binned output probabilities used for boson-sampling verification [61, 62, 63, 64].
Finally, we apply our method to near-Clifford circuits [65, 66] and show that we can efficiently estimate the expectation values of observables and any (marginal) output probabilities for these circuits within an additive-error.
Approximation of expectation values.— Let $|\psi\rangle=\otimes_{i=1}^{M}|\psi_{i}\rangle$ be a product input state in an $M$-mode system, and $\hat{U}$ be a quantum circuit applied to the input state. For the output state of the circuit, $\hat{U}|\psi\rangle$, we consider the expectation value of a product operator $\hat{O}=\otimes_{i=1}^{M}\hat{O}_{i}$ (see Fig. 1(a)):
$\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle.$ (1)
Throughout this work, input states $|\psi\rangle$ and operators $\hat{O}$ are assumed in a product form unless stated otherwise, which is practically relevant in various applications. When an operator acts nontrivially only on a subsystem $A$ and trivially on the rest of the system $B$, we write the operator as $\hat{O}=\hat{O}_{A}\otimes\hat{\mathbb{1}}_{B}=\bigotimes_{i\in A}\hat{O}_{i}\otimes\bigotimes_{i\in B}\hat{\mathbb{1}}_{i}$ and define the reduced density matrix $\hat{\rho}_{A}\equiv\mbox{Tr}_{B}[\hat{U}|\psi\rangle\langle\psi|\hat{U}^{\dagger}]$.
We first focus on a linear-optical circuit $\hat{U}$. Our first main result is a classical algorithm that approximates the expectation values of operators in a linear-optical circuit:
Theorem 1 (Expectation-value approximation in linear-optical circuit).
Consider an $M$-mode linear-optical circuit $\hat{U}$ and an operator $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ can be approximated within additive-error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log(1/\delta)/\epsilon^{2})$, where $\|\hat{O}_{A}\|_{2}^{2}=\mbox{Tr}[\hat{O}_{A}^{\dagger}\hat{O}_{A}]$.
Proof sketch.
(See SM Sec. S3 [40] for the full proof.) Assume that an operator $\hat{O}$ acts nontrivially on an $l$-mode subsystem $A$. Using the displacement operator expansion of the operator, $\hat{O}_{A}=\pi^{-l}\int d^{2l}\alpha\chi_{\hat{O}_{A}}(\bm{\alpha})\hat{D}^{\dagger}(\bm{\alpha})$ [67, 68, 69, 40], where $\hat{D}(\bm{\alpha})$ is a displacement operator and $\chi_{\hat{Q}}(\bm{\alpha})\equiv\mbox{Tr}[\hat{Q}\hat{D}(\bm{\alpha})]$ is the characteristic function of the operator $\hat{Q}$, we rewrite the expectation value as
$\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ $=\int d^{2l}\alpha\frac{|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}}{\pi^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})}{\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})},$ (2)
where $q(\bm{\alpha})\equiv|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}/(\pi^{l}\|\hat{O}_{A}\|_{2}^{2})$ is normalized to 1 and thus a proper probability distribution of $\bm{\alpha}$. Thus, the expression implies that by sampling $\bm{\alpha}$ from $q(\bm{\alpha})$ and averaging the random variable $X(\bm{\alpha})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})/\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})$ over the samples, we obtain an estimate of the expectation value. Thanks to the tensor-product structure of the state $\hat{\rho}_{A}$ and the operator $\hat{O}_{A}$, this process can be performed efficiently [70, 71, 72].
We then show that the median-of-means estimator [73] yields an estimate with additive-error $\epsilon$ with probability $1-\delta$ when $O(\|\hat{O}_{A}\|_{2}^{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log\delta^{-1}/\epsilon^{2})$ number of samples is given, by showing that the variance of the random variable $X(\bm{\alpha})$ is upper-bounded by $\|\hat{O}_{A}\|_{2}^{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]$. Note that the median-of-means estimator for complex random variables can be constructed by combining the real and imaginary parts estimators [40]. Since it takes $O(M^{2})$ time for computing random variable $X(\bm{\alpha})$, the total running time is $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log(1/\delta)/\epsilon^{2})$. ∎
Therefore, we achieve an efficient additive estimation of the expectation value in polynomial time when $\lVert\hat{O}_{A}\rVert^{2}_{2}=O(\text{poly}(M))$. Such a condition holds whenever $\hat{O}_{A}$ has $O(\text{poly}(M))$ rank with polynomially large singular values. We present two important classes of observables that satisfy the condition.
A first important class arises when $\hat{O}_{A}$ is local observables, directly relevant to photonic VQAs implemented by linear-optical circuits [45, 46, 47, 48, 49, 74, 50]. In these schemes, the physically relevant target Hamiltonian decomposes into few-mode local terms,
$\hat{H}=\sum_{i}\alpha_{i}\hat{H}_{i}+\sum_{i<j}\beta_{ij}\hat{H}_{ij}+\cdots,$ (3)
and each $\hat{H}_{i},\hat{H}_{ij},\dots$ acts on a small number of modes with bounded Hilbert–Schmidt norm as $O(1)$. Our algorithm efficiently estimates the contribution of each local term, and the total energy follows by summation. The same bounded-norm observables also appear as the primary measurement targets in other NISQ paradigms, such as analog and digital quantum simulators [75, 76, 77, 78, 79, 80, 81], where physical quantities, such as magnetizations, correlations, or energy densities, are expressed as few-body expectation values. Hence, the condition $\|\hat{O}_{A}\|_{2}^{2}=O(\mathrm{poly}(M))$ precisely characterizes the physically relevant observables not only in photonic VQAs but broadly across NISQ devices, delineating a classically simulable regime for expectation-value estimation in linear optics (See SM Sec. S2 for more details [40].).
A second important class consists of observables where $\hat{O}_{A}$ is a rank-one projector of product form. This case is directly tied to projective measurements on the output of a linear-optical circuit. A notable example is a photon-number projector in boson sampling [2]. In this case, Theorem 1 implies that the (marginal) output probabilities of boson sampling with arbitrary input states can be efficiently approximated. To the best of our knowledge, such a result has only been known for the Fock state input [2, 62, 39]. Remarkably, this finding also enables an efficient simulation of boson sampling with arbitrary product input if polynomially many outcomes dominate the output probabilities [57, 51, 82]. (Note that sparsity of the output distribution can be determined. See SM Sec. S9 for more details [40].)
Corollary 1 (Sparse boson sampling).
Boson sampling with an arbitrary product input state can be efficiently simulated if its output probability distribution is polynomially sparse.
In addition, note that quantum sampling with a peaked output probability distribution, i.e., one that has an outcome with an inverse-polynomially large probability, has drawn attention for a verifiable quantum advantage of sampling [55, 56] because the peak can serve as a verification. However, our result, together with the algorithm in Refs. [57, 51], indicates that the peak of the output probability distribution in boson sampling with any product input can be efficiently identified using classical computers. We emphasize that Corollary 1 is valid for a general case regardless of phase space negativity [83, 54] or stellar rank [84] of the input state, resources for quantum advantages.
Approximation of transition amplitudes.— We now present another classical algorithm approximating a transition amplitude $\langle\phi|\hat{U}|\psi\rangle$ (See Fig. 1 (b)). Note that when $|\phi\rangle$ and $|\psi\rangle$ are single-photon Fock states, the amplitude becomes a permanent [2], which can be efficiently approximated by Gurvits’ algorithm [58]. In contrast, our algorithm works for arbitrary product inputs.
Theorem 2 (Transition amplitude approximation in linear-optical circuit).
The transition amplitude $\langle\phi|\hat{U}|\psi\rangle$ of a linear-optical circuit $\hat{U}$ with arbitrary states $|\psi\rangle$ and $|\phi\rangle$ can be approximated within additive-error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$.
Proof Sketch.
We rewrite the amplitude by exploiting the completeness relation $\hat{\mathbb{1}}_{1:M}=\int d^{2M}\alpha|\bm{\alpha}\rangle\langle\bm{\alpha}|/\pi^{M}$:
$\langle\phi|\hat{U}|\psi\rangle$ $=\int d^{2M}\alpha\frac{\langle\phi|\hat{U}|\bm{\alpha}\rangle}{\langle\bm{\alpha}|\psi\rangle^{*}}\frac{|\langle\bm{\alpha}|\psi\rangle|^{2}}{\pi^{M}},$ (4)
where $|\bm{\alpha}\rangle$ represents a coherent state. Thus, if we sample $\bm{\alpha}$ from $|\langle\bm{\alpha}|\psi\rangle|^{2}/\pi^{M}$, which is a proper probability distribution (Husimi Q-distribution of the state $|\psi\rangle$ [85]) and easy to sample from, and average $\langle\phi|\hat{U}|\bm{\alpha}\rangle/\langle\bm{\alpha}|\psi\rangle^{*}$ over the samples, we obtain an estimate of the amplitude. We prove that this procedure is efficient and gives an estimate with $\epsilon$ additive-error with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$ in SM Sec. S4 [40]. ∎
Note that the original Gurvits’ algorithm applies only when $|\psi\rangle=|\bm{m}\rangle$, $|\phi\rangle=|\bm{n}\rangle$ with $\bm{n},\bm{m}\in\{0,1\}^{M}$, and its variants [86, 39] apply when $\bm{m}\in\{0,1\}^{M}$ or $\bm{n}=\bm{m}$, corresponding to approximating the permanent of matrices with repeated rows and columns in a certain pattern. Notably, our algorithm not only extends to computing the permanent of matrices with repeated rows and columns in a general pattern, but also introduces a new approach to approximating the permanent using a quantum-optics-inspired method.
Our algorithm can also approximate the hafnian [59], a generalization of the permanent, by using the fact that the hafnian of any $M\times M$ complex symmetric matrix $R$ ($M$ is even) with the largest singular value $\lambda_{\text{max}}<1$ can be expressed as a transition amplitude of an $M$-mode Gaussian boson sampling circuit [3, 60]:
$\text{haf}(R)=\frac{1}{Z^{1/2}}\langle\bm{1}|\hat{U}|\bm{r}\rangle,$ (5)
[…]
Corollary 2 (Hafnian approximation).
The hafnian of an $M\times M$ complex symmetric matrix $R$ can be approximated within additive-error $\epsilon\norm{R}^{M/2}$ with probability $1-\delta$ in $O(M^{3}\log\delta^{-1}/\epsilon^{2})$ time.
A detailed proof is provided in SM Sec. S5 [40]. By a similar method, the loop hafnian of an arbitrary complex symmetric matrix [60] can also be efficiently approximated.
It is worth emphasizing that Theorem 2 can also be used to approximate the overlap of two quantum states generated by linear-optical circuits. Such a quantity has attracted attention for machine learning applications of boson sampling, such as kernel method [33, 34, 35]. More specifically, boson samplers are used to estimate the overlap of two quantum states corresponding to two different data vectors $\bm{x},\bm{y}$ as $|\langle\psi|\hat{U}^{\dagger}(\bm{x})\hat{U}(\bm{y})|\psi\rangle|^{2}$, where $\hat{U}(\bm{x})$ corresponds to the linear-optical circuits whose configuration depend on the data vector $\bm{x}$. Thus, our classical method can replace boson samplers for this purpose.
Phase shifter.— We now consider a problem of approximating the expectation value of a phase shifter $e^{i\hat{\bm{n}}\cdot\bm{\phi}}$ of a linear-optical circuit [87]
$\langle\psi|\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}|\psi\rangle,$ (6)
where $\hat{\bm{n}}\equiv(\hat{n}_{1},\dots,\hat{n}_{M})$ and $\bm{\phi}\in\mathbb{R}^{M}$ represents the phase vector. The expectation value of a phase shifter has various applications such as the MVS problem [31, 39] and approximating binned probability, proposed for verifying boson sampling experiments [61, 62, 63, 64].
In particular, the MVS problem [31] has been considered as a potential application of boson sampling, which is essentially equivalent to computing grouped output probabilities $p(\bm{m})$ of boson sampling (See SM Sec. S4 [40] for more details):
$G(\Omega)\equiv\sum_{\bm{m}=\bm{0}}^{\infty}p(\bm{m})\delta(\Omega-\bm{\omega}\cdot\bm{m}),$ (7)
where $\bm{\omega}\in\mathbb{Z}^{M}_{\geq 0}$ and $\Omega=\{0,\dots,\Omega_{\text{max}}\}$. In Ref. [39], however, the authors observe that the Fourier components of Eq. (7) are expressed as the expectation value of a phase shifter as Eq. (6). Therefore, if Eq. (6) is additively approximated, the grouped probabilities Eq. (7) are reproduced within an additive-error through the inverse Fourier transform [40].
While Ref. [39] proves that this is the case when $|\psi\rangle$ is a product Gaussian state or a Fock state, leading to dequantization, for more general input states, like a product squeezed Fock state, a quantum advantage may still exist since no classical algorithm is known yet. This possibility is particularly important because such an extension has recently been experimentally implemented [38]. Notably, by setting $|\phi\rangle=|\psi\rangle$ and $\hat{U}\to\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ in Theorem 2, our classical algorithm can approximate Eq. (6) for any product input state:
Corollary 3 (Phase-shifter expectation value).
Phase-shifter expectation value of a linear-optical circuit can be efficiently approximated for an arbitrary product input state.
Therefore, our algorithm enables efficient approximation of the grouped probabilities in Eq. (7) for arbitrary product input states and thus solves a generalized MVS problem for arbitrary product input states, an open problem posed in Ref. [39].
Extending to qubit circuits.— Finally, we show that our classical algorithm in Theorem 1 can be extended to qubit system, especially near-Clifford circuits $\hat{U}_{\text{NC}}$ [65, 66], i.e., Clifford circuit $\hat{U}_{\text{C}}$+logarithmic number of $T$ gates, with arbitrary product input states. It is known that both exact and approximate classical simulations of such circuits are computationally hard, assuming the non-collapse of the polynomial hierarchy [88]. Consequently, to investigate routes toward practical quantum advantage, our method provides a direct additive-error estimation scheme for the expectation value of an observable, $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}\hat{U}_{\text{NC}}|\psi\rangle$ that does not rely on sampling from the output distribution, thereby bypassing the main bottleneck that underlies the known sampling hardness results.
Crucial properties used in Theorem 1 are that a displacement operator is transformed into another displacement operator through linear-optical circuits and that any operator can be expanded in terms of displacement operators. Similar properties hold for Clifford circuits where any operators can be decomposed using Pauli operators and a Pauli operator transforms another Pauli operator under Clifford circuits. Thus, we obtain a similar result:
Theorem 3 (Expectation value approximation in near-Clifford circuit).
Consider an $n$-qubit near-Clifford circuit $\hat{U}_{\text{NC}}$ with depth $O(\text{poly}(n))$ and an operators $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}\hat{U}_{\text{NC}}|\psi\rangle$ can be approximated within additive-error $\epsilon$ with probability $1-\delta$ in running time $O(\text{poly}(n)\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log(1/\delta)/\epsilon^{2})$.
Proof Sketch.
The proof is similar to that of Theorem 1 and provided in SM Sec. S6 [40]. Let us first focus on Clifford circuits. Using the Pauli operator expansion, $\hat{O}_{A}=2^{-l}\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\chi_{\hat{O}_{A}}(\bm{a})\hat{P}_{\bm{a}}$, where $\chi_{\hat{Q}}(\bm{a})\equiv\mbox{Tr}[\hat{Q}\hat{P}_{\bm{a}}]$ and $\hat{P}_{\bm{a}}\in\{I,X,Y,Z\}^{l}$ is a Pauli operator, the expectation value of $\hat{O}=\hat{O}_{A}\otimes\hat{\mathbb{1}}_{B}$ can be written as
$\langle\psi|\hat{U}^{\dagger}_{\text{C}}\hat{O}\hat{U}_{\text{C}}|\psi=\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\frac{|\chi_{\hat{O}_{A}}(\bm{a})|^{2}}{2^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})}{\chi^{*}_{\hat{O}_{A}}(\bm{a})}.$ (8)
Thus, by sampling $\bm{a}$ from $q(\bm{a})\equiv|\chi_{\hat{O}_{A}}(\bm{a})|^{2}/(2^{l}\|\hat{O}_{A}\|_{2}^{2})$ and averaging the random variable $X(\bm{a})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})/\chi^{*}_{\hat{O}_{A}}(\bm{a})$ over the samples, we obtain an estimate of the expectation value. By showing that the variance of the random variable $X(\bm{a})$ is bounded by $\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]$, the median-of-means estimator gives the desired error presented in the theorem.
Furthermore, our method is still valid for the additional logarithmic number of $T$-gates because a single Pauli operator is transformed into the sum of at most two Pauli operators after passing through each $T$-gate, that is, $\hat{T}^{\dagger}\hat{P}_{0}\hat{T}=\hat{P}_{0},~\hat{T}^{\dagger}\hat{P}_{1}\hat{T}=(\hat{P}_{1}+\hat{P}_{2})/\sqrt{2},~\hat{T}^{\dagger}\hat{P}_{2}\hat{T}=(\hat{P}_{2}-\hat{P}_{1})/\sqrt{2},~\hat{T}^{\dagger}\hat{P}_{3}\hat{T}=\hat{P}_{3}$. If there are $t$ of $T$-gates in a near-Clifford circuit $\hat{U}_{\text{NC}}$, we have at most $n2^{t}$ Pauli operators to track the transformations. Therefore, if $t=O(\log n)$, we can handle all the relevant Pauli operators in $O(\text{poly}(n))$ time. A detailed proof is given in SM Sec. S7 [40]. ∎
Theorem 3 implies that the expectation value of observables $\hat{O}$ with $\lVert\hat{O}\rVert_{2}^{2}=O(\text{poly}(n))$ of a near-Clifford circuit with an arbitrary product input state can be efficiently approximated. Especially for rank-1 projective measurements, this result implies that the magic resources of both the input state and measurement do not increase the complexity of additive-error-approximation. Note that our result far extends the previous results: the expectation values of a near-Clifford circuit with stabilizer input or for Pauli observables [65, 66], and output probabilities in the computational basis [89, 66, 82]. (See SM Sec. S10 for more details [40].)
Discussion.— We presented a classical algorithm for estimating expectation values of observables in linear-optical circuits. Using this, we showed that one can efficiently approximate (marginal) output probabilities of boson sampling with arbitrary product input states, which implies that the simulation of sparse boson sampling with arbitrary product input states is easy. We also provided an efficient classical algorithm for the general transition amplitude in a linear-optical circuit, which allows us to approximate the hafnian of any complex symmetric matrix. Furthermore, using this classical algorithm, we solved an open problem in Ref. [39], namely, a generalized MVS problem for arbitrary product input states.
A natural open question is to extend our method to more complicated optical circuits with nonlinear effects such as nonlinear gates or postselection. Since such cases ultimately enable universal quantum computation [90, 1], we do not expect our method to work for the most general cases. Nonetheless, it suggests a phase transition from easiness to hardness when introducing nonlinear effects; thus, an interesting open question is to determine the exact transition point.
Also, we have focused on computing grouped probabilities of boson sampling with a certain pattern in the MVS problem, which is defined by the linear relation $\Omega-\bm{\omega}\cdot\bm{m}$ in Eq. (7). However, our method does not straightforwardly generalize to a more complicated pattern of groups, which still opens a possibility of quantum advantage. Thus, the complexity of approximating general grouped boson sampling output probabilities is another open question.
Acknowledgements.
Acknowledgments.— This work was supported by the National Research Foundation of Korea Grants (No. RS-2024-00431768 and No. RS-2025-00515456) funded by the Korean government (Ministry of Science and ICT (MSIT)), the Institute of Information & Communications Technology Planning & Evaluation (IITP) Grants funded by the Korea government (MSIT) (No. IITP-2025-RS-2025-02283189 and IITP-2025-RS-2025-02263264), and by Creation of the quantum information science R&D ecosystem (based on human resources) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT (MSIT)) (RS-2023-00256050).